Estimate the magnetic field 1 m away from a lightning strike
Answer #1

lercure l m 27d 3 叉下大 二, 60X10
Answer #2

Lightning striking the ground has a range of current of 200 kA

So B* 2*pi*r = mu * I

So B = 200 * 10^3 * 2 * 10^-7 /1 = 4* 10^-2 = 0.04 T


answered by: Thebesta